# A Proof of the Weak Gravity Conjecture via Leading-Order Quantum Gravity Effects

### v2

@@ V Concluding remarks
In this paper, we have explored the effective action obtained by integrating out massless particles at the one-loop level. We confirm that the loop effects of quantum gravity can naturally induce super-extremality ($q>M$) in charged black holes Arkani-Hamed:2021ajd . It suggests that quantum gravity plays a crucial role in facilitating the decay of extremal black holes into more fundamental components. These effects thus have far-reaching implications for our understanding of the ultimate evolution of black holes and the universe.
We emphasize that these results are independent of specific UV theories of quantum gravity. Instead, they depend solely on the low-energy particle spectrum of nature, which endows them with predictive power. Specifically, by using the particle spectrum of the Standard Model with a minimally coupled Higgs and taking $\mu_{*}$ as the Planck energy, we predict that for a black hole with a charge of approximately $10^{-16}C$ (the typical charge carried by a dust particle), the quantity $\frac{q}{M}-1\approx 5.2\times 10^{-5}$.
Finally, we have provided a comprehensive thermodynamic analysis of both extremal and non-extremal RN black holes, which has further interesting implications. For extremal RN black holes, it unexpectedly reveals a connection between super-extremality and the logarithmic entropy corrections, which were previously viewed as two independent research directions. For non-extremal RN black holes, we also demonstrate a relationship between the shift in the charge-to-mass ratio and the thermodynamic stability properties of the black holes. Both of these aspects merit further investigation.

### v3

@@ V Concluding remarks
In this paper, we have investigated the effective action obtained from integrating out massless particles at the one-loop level. Our results confirm that the loop effects of quantum gravity can naturally induce super-extremality ($q>M$) in charged black holes Charles:2019qqt ; Arkani-Hamed:2021ajd . It suggests that quantum gravity plays a crucial role in facilitating the decay of extremal black holes into more fundamental components. These effects thus have far-reaching implications for our understanding of the ultimate evolution of black holes and the universe. We emphasize that the results are model-independent, as they depend solely on the low-energy particle spectrum of nature and are insensitive to the UV details of quantum gravity theories.
We have provided the thermodynamic analysis for both extremal and non-extremal RN black holes, which has interesting implications that warrant further investigation. For extremal RN black holes, our work unexpectedly reveals a connection between the super-extremality of WGC and the logarithmic entropy corrections, which were previously regarded as separate research directions. For non-extremal RN black holes, we establish a relationship between the shift in the charge-to-mass ratio and the thermodynamic stability properties of the black holes. A general discussion on the correlation between thermodynamic stability and super-extremality can be found in Cheung:2018cwt , where it was formalized based on tree-level higher-derivative terms. Our results suggest that this correlation can also be extended to the one-loop effects and non-extremal cases.